# FreqKV: Frequency Domain Key-Value Compression for Efficient Context Window Extension

### v2

@@ 5 Experiments
5.1 Implementation
We conduct experiments on long context language modeling and understanding tasks with LLaMA-2-7b (Touvron et al.,, 2023) and LLaMA-3-8b (AI@Meta,, 2024). We maintain $S=4$ sink tokens uncompressed. The retaining ratio $\gamma$ in compression is set to 0.5. The preset context window size of LLaMA-2 is 4096, which is also the maximum KV cache size $N$. Therefore, the retaining size during each compression is $L=\gamma\cdot(N-S)=2046$. As long as the cache size reaches its capacity of 4096, the 4092 states since the $5$-th state in the cache will be compressed into 2046 states. As for LLaMA-3, its context size is 8192. It is equipped with GQA (Grouped-Query Attention), which means it has a lower proportion of parameters for attention modules than LLaMA-2-Base/Chat-7b (Multi-Head Attention, MHA).
Minimal training is introduced to adapt models to this frequency-domain compression method. For long context language modeling, we fine-tune LLaMA-2-base on the pre-training dataset RedPajama (Computer,, 2023), extending the context window from 4K to 8K, 16K, and 32K. Perplexity (PPL) evaluation is conducted on PG-19 (Rae et al.,, 2019) and Proof-pile (Azerbayev et al.,, 2022). For long context understanding, the instruction following dataset LongAlpaca (Chen et al.,, 2024) is used for the supervised fine-tuning (SFT) of LLaMA-2-chat and LLaMA-3-instruct. The context window is extended from 4K to 8K, and from 8K to 16K respectively. Models are evaluated on LongBench (Bai et al.,, 2024) and Needle-in-a-Haystack (gkamradt,, 2023). Details of training settings are summarized in Appendix C.
Table 1: Perplexity evaluation on the test sets of PG-19 and Proof-pile. The superscript “*” means that we reproduce LoCoCo following their official code for evaluation. The results of full fine-tuning and LongLoRA are reported from (Chen et al.,, 2024).
Training Length Method Inference Cache Evaluation Context Length
@@ 5.1 Implementation
PG-19
8192 Full FT Full 7.55 7.21 6.98 - -
@@
16384 LongLoRA Full 7.65 7.28 7.02 6.86 -
@@
Proof-pile
8192 Full FT Full 3.14 2.85 2.66 - -
LongLoRA Full 3.20 2.91 2.72 - -
LoCoCo* Compressed 3.40 3.20 2.88 - -
FreqKV Compressed 3.15 2.88 2.79 2.76 2.75
16384 LongLoRA Full 3.17 2.87 2.66 2.51 -
FreqKV Compressed 3.15 2.88 2.78 2.74 2.73
32768 LongLoRA Full 3.35 3.01 2.78 2.61 2.50
FreqKV Compressed 3.16 2.89 2.78 2.74 2.72
@@ 5.2 Long Context Language Modeling
We use FreqKV to train LLaMA-2-base-7b on RedPajama with lengths of 8192, 16384, and 32768. Perplexity is measured on test sets of the book corpus dataset PG-19 and the Arxiv math dataset Proof-pile with the evaluation sliding window of 256. We compare our method with other baselines including full fine-tuning (Full FT), LongLoRA (Chen et al.,, 2024), and LoCoCo (Cai et al., 024a, ). While Full FT and LongLoRA leverage full KV cache during inference, LoCoCo, and our FreqKV use compressed cache.
PPL scores on different evaluation context lengths are reported in Table 1. While both of LongLoRA and LoCoCo sacrifice performance within the original context length (2K and 4K), FreqKV almost has no performance degradation. Even with the training length of 8K, FreqKV performs comparably to LongLoRA with the training length of 32K. It also outperforms LoCoCo on the extended context length as well as the shorter lengths.
It can be noticed that FreqKV falls behind LongLoRA on Proof-pile at longer context lengths. Proof-pile consists of a diverse set of mathematical texts, making it particularly challenging for models using compressed cache to predict the next token. In contrast, LongLoRA retains full KV cache during inference, which naturally benefits next-token prediction in such structured domains. However, the performance of FreqKV does not deteriorate as the context length increases. Our compressed cache method even surpasses Full FT and LongLoRA on PG-19.
Furthermore, we evaluate the performance of FreqKV at longer lengths as shown in Figure 5. With the training length of 8K, it effectively extends the context window to 64K, which is 16 times larger than the original size.
5.3 Long Context Understanding
To further validate the performance on downstream tasks, we use FreqKV to extend the context window of LLaMA-2-chat-7b from 4K to 8K and LLaMA-3-instruct-8b from 8K to 16K. Models are evaluated on the long context understanding benchmark LongBench and Needle-in-a-Haystack.
LongBench.
Scores on the 5 categories of LongBench are reported in Table 2. FreqKV consistently improves the performance of LLaMA-2-chat and LLaMA-3-instruct across a majority of these tasks. We compare our method with different KV compression strategies, including LM-Infinite (Han et al.,, 2024), LongHeads (Lu et al.,, 2024), FastKV (Jo et al.,, 2025), SnapKV (Li et al.,, 2024) and PyramidKV (Cai. et al., 024b, ). Most baselines use LLaMA-2-chat-7b as their backbone. The retaining ratio is set to 0.5. FreqKV achieves SOTA (state-of-the-art) on most of the long context understanding tasks and obtains the highest average score.
Needle-in-a-Haystack.
The Needle-in-a-Haystack results are illustrated in Figure 3. FreqKV performs well in extending the context window of LLaMA-2 and LLaMA-3. In contrast, PyramidKV fails beyond the original window boundary. This is because it compresses the long prompt after the first decoding step. Out-of-bound position embeddings are used and it prunes tokens after calculating attention weights. Consequently, those unseen positions lead to the performance collapse.
Table 2: Scores of different KV compression methods on LongBench. The superscript “*” means that we reproduce FastKV, SnapKV and PyramidKV following their official codes for evaluation. The results of LM-Infinite and LongHeads are reported from Lu et al., (2024).
Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Avg.
@@ 5.3 Long Context Understanding
LLaMA2-chat 18.7 19.2 36.8 25.4 32.8 9.4 27.3 20.8 25.8 61.5 77.8 40.7 52.4 43.8 35.17
+LM-Infinite 0.00 18.57 25.33 27.34 31.96 7.76 11.30 2.99 8.72 32.50 29.22 13.82 34.19 24.55 19.16
+LongHeads 11.61 22.98 23.76 31.28 24.10 8.87 25.36 20.24 16.18 50.67 79.98 36.74 53.85 44.22 32.13
+FastKV* 18.25 22.46 35.83 28.84 31.34 8.49 24.4 20.96 25.8 60.0 82.67 40.64 57.43 51.52 36.33
+SnapKV* 17.91 22.9 35.3 26.03 28.21 8.77 24.99 20.94 25.92 64.0 83.36 41.0 60.83 55.14 36.81
+PyramidKV* 17.36 23.58 34.89 26.28 28.39 9.17 24.95 20.9 26.04 64.0 83.36 40.92 60.68 55.17 36.84
+FreqKV 20.41 21.05 31.2 34.54 34.78 14.47 25.51 21.81 26.9 56.0 84.09 41.53 58.99 58.65 37.85
LLaMA3-instruct 22.52 31.83 41.04 44.14 36.68 24.05 28.87 23.25 26.46 75.5 90.23 42.09 56.7 51.84 42.51
+PyramidKV* 22.27 31.93 41.03 44.79 36.56 23.86 28.77 23.14 26.62 74.5 90.25 42.53 59.11 53.33 42.76
+FreqKV 23.14 41.6 43.62 46.48 40.54 20.2 29.99 23.67 28.58 69.0 88.95 41.47 60.33 51.11 43.48
(a) Results of PyramidKV on LLaMA-2. The average accuracy is 46.7%.
(b) Results of Dropping on LLaMA-2. The average accuracy is 75.5%.
(c) Results of FreqKV on LLaMA-2. The average accuracy is 84.2%.
(d) Results of FreqKV on LLaMA-3. The average accuracy is 90.7%.
@@
Figure 3: The Needle-in-a-Haystack results, with the x-axis representing the document length ranging from 1K to 8K tokens for LLaMA-2-chat-7b and 1K to 16K tokens for LLaMA-3-instruct-8b, and the y-axis showing the position of the “needle” within the document.
To demonstrate the effectiveness of our compression operation, a variant method Dropping is also implemented for comparison as in Figure 3b. Instead of compressing in the frequency domain, Dropping retains the most recent tokens like StreamingLLM (Xiao et al.,, 2024). It shares the same training settings as FreqKV. While Dropping suffers from the permanent loss of earlier tokens, FreqKV could retrieve information from a specific sentence (the “needle”) hidden within a document (the “haystack”) and achieve a better performance.
@@ 6 Analysis
Figure 4: Perplexity evaluation on the validation set of PG-19 with different retaining ratios.
6.1 Retaining Ratio
We train LLaMA-2-7b on RedPajama (Computer,, 2023) with different retaining ratios. The training length is 8K. Models are evaluated on the validation set of PG-19 with the evaluation length ranging from 2K to 64K.
The performance of FreqKV with different retaining ratios is presented in Figure 5. FreqKV performs better with more frequency components retained. Employing an iterative compression manner, less compressions are performed on the more recent tokens by FreqKV, which benefits long context language modeling. It consistently achieves low perplexity as the evaluation context length grows from 2K to 64K, far exceeding the training length.
6.2 Decoding Latency
Table 3: The decoding latency (seconds) of FreqKV and Dropping on LLaMA-2-7b with the increasing sequence length. The difference between FreqKV and Dropping shows the compression latency introduced by FreqKV.
Models 4K 8K 16K 24K 32K
Compression Times 0 3 7 11 15
Dropping 123.78$\pm$1.46 253.32$\pm$5.76 506.60$\pm$5.11 760.46$\pm$7.15 1022.50$\pm$18.83
FreqKV 124.34$\pm$1.80 252.75$\pm$3.55 507.61$\pm$5.44 764.82$\pm$9.72 1029.00$\pm$19.97
FreqKV performs compressions on-the-fly during decoding. We compare the decoding time required by LLaMA-2-7b with the full cache and our FreqKV when the sequence length increases. As shown in Figure 5, the decoding time starts to diverge at the length of 4K. While the full cache utilization leads to a quadratic growth in decoding time, the decoding time of FreqKV increases approximately linearly with a negligible time spent on compression, showcasing its efficiency.
Furthermore, we measure the decoding latency of FreqKV and Dropping with the increasing length. Dropping removes the earliest tokens instead of compression in the frequency domain. The difference in the decoding time between the two methods shows the latency of compression. The sink size is set to 4 and the retaining ratio is 0.5. Table 3 reports the average decoding time (seconds) over five runs, along with standard deviations. It can be observed that the additional latency (6.5s) introduced by FreqKV accounts for only 0.64% of the total decoding time at 32K tokens. Further analysis for the compression overhead is provided in Appendix D.

### v3

@@ 4 Experiments
4.1 Experimental Settings
We employ FreqKV on LLaMA-2-7B/13B (Touvron et al., 2023) and LLaMA-3-8B (AI@Meta, 2024). Unless otherwise specified, the retaining ratio $\gamma$ is set to 0.5, and $S=4$ sink tokens are kept uncompressed. For LLaMA-2, the preset context window size is 4096, which also defines the maximum KV cache size $N$. Accordingly, during each compression step, the NS=4092N-S=4092 states since the $5$-th state in the cache will be compressed into $L=\gamma\cdot(N-S)=2046$ states. As for LLaMA-3, the native context window size is 8192. It is equipped with GQA (Grouped-Query Attention), which means it has a lower proportion of parameters for attention modules than LLaMA-2 (Multi-Head Attention, MHA). Minimal training is introduced to adapt models to this frequency-domain compression method, following LongLoRA (Chen et al., 2024). Details of training settings are summarized in Appendix D.
We assess FreqKV on long context language modeling and generation tasks for decoding performance, and long context understanding tasks for prefilling performance. For long context language modeling, we fine-tune LLaMA-2-base on the pre-training dataset RedPajama (Computer, 2023), extending the context window from 4K to 32K. Perplexity (PPL) evaluation is conducted on PG-19 (Rae et al., 2019). For long context generation and understanding, the instruction following dataset LongAlpaca (Chen et al., 2024) is used for the supervised fine-tuning (SFT) of LLaMA-2-chat and LLaMA-3-instruct. The context window is extended from 4K to 8K, and from 8K to 16K respectively. Models are evaluated on LongGenBench (Liu et al., 2024), LongBench (Bai et al., 2024), RULER (Hsieh et al., 2024), and Needle-in-a-Haystack (gkamradt, 2023).
Table 2: Perplexity evaluation on the test set of PG-19. The superscript “*” means that we reproduce LoCoCo following their official code for evaluation. The results of full fine-tuning and LongLoRA are reported from LongLoRA (Chen et al., 2024). Full FT and LoCoCo encounter the OOM (Out-of-Memory) issue with the training length of 16K.
Size Training Length Method Inference Cache Evaluation Context Length
@@ 4.1 Experimental Settings
7B 8192 Full FT Full 7.55 7.21 6.98 - -
LoCoCo* Compressed 8.15 8.08 7.27 - -
@@
16384 Full FT Full OOM during Training
LoCoCo Compressed OOM during Training
LongLoRA Full 7.65 7.28 7.02 6.86 -
@@
13B 8192 Full FT Full 6.95 6.60 6.43 - -
LongLoRA Full 7.03 6.73 6.58 - -
FreqKV Compressed 6.82 6.52 6.44 6.42 6.41
@@ 4.2 Long Context Language Modeling
We use FreqKV to train LLaMA-2-7B and LLaMA-2-13B on RedPajama. Perplexity is measured on the test set of the book corpus dataset PG-19 with the evaluation sliding window of 256. We compare our method with other baselines, including full fine-tuning (Full FT), LongLoRA, and LoCoCo (Cai et al., 2024). While Full FT and LongLoRA leverage full KV cache during inference, LoCoCo and FreqKV use compressed cache.
PPL scores on different evaluation context lengths are reported in Table 2. More Evaluation results on Proof-pile (Azerbayev et al., 2022) are presented in Appendix E. While both LongLoRA and LoCoCo sacrifice performance within the original context length (2K and 4K), FreqKV matches or even surpasses Full FT. With the training length of 8K, FreqKV performs comparably to LongLoRA trained at 32K and consistently outperforms LoCoCo in terms of extended context length as well as the shorter lengths. Additional ablation studies on retained frequency components, retaining ratio, and sink size are reported in Appendix F.
Table 3: Scores of different KV compression methods on LongBench. The superscript “*” means that results of CaM and KVMerger are reported from KVMerger (Wang et al., 2024). We reproduce the other baselines following their official codes for evaluation. The second-best scores are also underlined in the table.
Retaining Ratio Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Avg.
@@
LLaMA-2-chat-7B
100% Full Cache 18.7 19.2 36.8 25.4 32.8 9.4 27.3 20.8 25.8 61.5 77.8 40.7 52.4 43.8 35.17
50% CaM* 17.08 20.00 33.98 - 30.69 - 24.46 - 24.66 63.5 82.17 - - - -
KVMerger* 18.50 20.04 36.89 - 32.99 - 25.31 - 26.29 64.0 83.62 - - - -
LaCache 16.27 16.55 31.55 32.62 26.22 7.72 22.84 20.34 25.16 65.0 83.28 38.46 47.97 43.41 34.10
D2O 20.32 19.16 33.09 30.62 26.76 9.65 23.72 20.54 25.42 63.0 84.23 41.22 58.61 52.77 36.37
GemFilter 15.22 17.76 32.92 27.29 26.76 8.66 16.17 16.29 15.95 57.5 84.8 31.96 46.54 45.14 31.64
FastKV 18.25 22.46 35.83 28.84 31.34 8.49 24.4 20.96 25.8 60.0 82.67 40.64 57.43 51.52 36.33
SnapKV 17.91 22.9 35.3 26.03 28.21 8.77 24.99 20.94 25.92 64.0 83.36 41.0 60.83 55.14 36.81
ThinK 15.57 21.76 37.98 30.21 31.55 8.06 26.24 20.98 26.1 65.0 80.08 40.97 60.2 54.26 37.07
PyramidKV 17.36 23.58 34.89 26.28 28.39 9.17 24.95 20.9 26.04 64.0 83.36 40.92 60.68 55.17 36.84
FreqKV 20.41 21.05 31.2 34.54 34.78 14.47 25.51 21.81 26.9 56.0 84.09 41.53 58.99 58.65 37.85
1% LaCache 4.37 12.72 6.34 14.77 14.91 2.23 11.18 7.39 9.94 16.75 15.13 6.42 7.56 7.09 9.77
D2O 15.53 12.61 17.64 29.0 21.71 8.49 16.56 16.05 16.63 52.5 81.03 20.2 31.88 33.47 26.66
FastKV 10.55 13.49 13.53 13.27 6.8 4.11 11.94 14.43 11.92 27.0 52.05 17.22 32.1 29.38 18.41
GemFilter 2.36 3.22 3.99 1.28 2.09 0.78 11.07 13.14 10.9 12.5 36.38 15.54 21.16 26.61 11.50
SnapKV 11.49 12.47 13.61 12.79 7.05 3.48 11.96 14.61 12.16 28.0 52.11 17.03 32.48 28.39 18.40
ThinK 8.38 14.5 14.71 20.13 22.21 4.94 11.97 16.44 11.63 18.0 19.59 9.46 21.87 17.09 15.07
PyramidKV 18.17 23.8 35.97 28.42 31.7 8.76 15.85 16.64 16.2 64.5 83.65 32.41 60.38 54.89 35.10
FreqKV 19.44 20.2 31.68 35.89 32.0 12.52 15.96 17.39 15.91 60.5 83.98 33.07 62.47 56.57 35.54
LLaMA-3-instruct-8B
100% Full Cache 22.52 31.83 41.04 44.14 36.68 24.05 28.87 23.25 26.46 75.5 90.23 42.09 56.7 51.84 42.51
50% LaCache 22.95 28.91 40.99 46.34 36.26 23.02 16.92 16.02 15.59 74.0 91.11 32.97 55.09 49.31 39.25
D2O 26.61 28.29 38.85 44.95 36.57 22.9 27.66 23.34 25.62 74.0 90.31 41.79 57.57 54.54 42.36
GemFilter 23.6 29.26 43.71 42.18 36.35 24.48 17.75 16.46 15.7 73.0 90.79 33.3 53.19 38.54 38.45
FastKV 23.88 29.67 40.75 44.66 34.02 23.07 16.86 16.45 15.61 74.0 90.31 42.02 59.21 53.3 39.63
SnapKV 22.99 31.89 41.11 43.92 36.42 23.38 17.14 16.45 15.53 74.5 90.23 33.72 59.03 53.97 40.02
ThinK 21.98 30.97 41.9 45.1 34.18 22.51 27.73 23.38 26.4 74.5 90.16 41.5 62.05 58.37 42.91
PyramidKV 22.27 31.93 41.03 44.79 36.56 23.86 28.77 23.14 26.62 74.5 90.25 42.53 59.11 53.33 42.76
FreqKV 23.14 41.6 43.62 46.48 40.54 20.2 29.99 23.67 28.58 69.0 88.95 41.47 60.33 51.11 43.48
1% LaCache 14.27 5.87 20.73 40.86 33.07 18.58 15.01 7.76 10.89 40.0 77.92 16.23 40.7 40.51 27.31
D2O 24.95 12.77 20.19 36.61 27.2 15.43 22.57 15.18 15.12 56.0 89.34 24.75 32.79 36.18 30.65
GemFilter 11.8 6.48 19.3 15.23 15.35 7.8 12.28 14.25 9.84 53.42 54.19 20.51 19.37 21.79 20.12
FastKV 19.57 12.61 31.28 42.59 29.44 20.71 13.96 15.32 12.86 61.0 87.79 29.55 56.74 54.18 34.83
SnapKV 18.29 14.25 31.87 38.03 26.82 20.03 14.29 14.57 13.08 60.5 88.52 29.95 55.89 53.36 34.25
ThinK 18.02 10.38 31.46 39.1 26.08 17.46 17.31 21.48 17.97 39.5 88.18 37.27 54.69 55.19 33.86
PyramidKV 22.79 31.91 41.24 44.2 36.67 22.53 17.29 16.52 15.57 74.5 90.23 33.3 58.95 53.12 39.92
FreqKV 22.86 39.7 36.44 44.29 39.59 22.57 29.1 22.92 28.35 65.0 89.0 41.9 62.17 55.42 42.81
4.3 Long Context Understanding
To further assess the performance on long-prefilling scenarios, we apply FreqKV to extend the context window of LLaMA-2-chat-7b from 4K to 8K and LLaMA-3-instruct-8b from 8K to 16K. Evaluation is conducted on three long context understanding benchmarks: LongBench, RULER, and Needle-in-a-Haystack.
LongBench.
Scores on the 5 categories of LongBench are reported in Table 3. FreqKV consistently improves the performance of LLaMA-2-chat and LLaMA-3-instruct across a majority of these tasks. We compare our method against eviction-based strategies GemFilter (Shi et al., 2024), SnapKV (Li et al., 2024), PyramidKV (Cai. et al., 2024b), LaCache (Shi et al., 2025), and FastKV (Jo et al., 2025), merging-based strategies CaM (Zhang et al., 2024), KVMerger (Wang et al., 2024), and D2O (Wan et al., 2025), and a channel compression method ThinK (Xu et al., 2025) that improves SnapKV. FreqKV achieves SOTA (state-of-the-art) on most of the long context understanding tasks and obtains the highest average score. In particular, it shows significant gains on question answering and summarization tasks such as HotpotQA, 2WikiMQA, and QMSum. Moreover, even with an extremely low retaining ratio ($\gamma=1\%$), where other methods degrade sharply, FreqKV remains stable and effective, highlighting its robustness under aggressive compression.
Table 4: Scores of different KV compression methods on RULER. D2O encounters the OOM issue with the evaluation length of 16K.
Evaluation Length Method CWE FWE QA VT NIAH Avg.
8K LLaMA-3 96.40 82.33 62.50 95.00 97.81 86.81
+D2O 93.20 81.00 62.00 88.80 47.53 74.51
+GemFilter 96.40 82.33 62.50 95.00 97.81 86.81
+FastKV 86.10 77.67 63.50 93.00 87.03 81.46
+SnapKV 96.00 80.33 62.50 93.60 92.28 84.94
+ThinK 95.7 78.67 62.00 96.8 94.91 85.62
+PyramidKV 96.40 82.33 62.50 95.00 97.81 86.81
  +FreqKV 91.80 84.67 64.50 100.00 97.44 87.68
16K +D2O - - - - - OOM
+GemFilter 0.00 0.00 0.50 0.00 0.00 0.10
+FastKV 0.00 0.00 0.00 0.00 0.00 0.00
+SnapKV 0.00 0.00 0.50 0.00 0.00 0.10
+ThinK 0.00 0.00 0.00 0.00 0.00 0.00
+PyramidKV 0.00 0.00 0.50 0.00 0.00 0.10
  +FreqKV 45.10 87.33 34.00 44.60 25.47 47.30
RULER.
Results on RULER are presented in Table 4. FreqKV consistently outperforms baselines across different tasks. Notably, baseline methods collapse entirely at 16K evaluation length, which exceeds the original context window of LLaMA-3-instruct (8K). This failure arises because they compress the long prompt after the first decoding step. Out-of-bound position embeddings are used when calculating attention, resulting in a performance breakdown. In contrast, FreqKV remains robust and demonstrates its effectiveness for context extension. FreqKV first conducts self-attention within the context window, and then compresses the KV states for out-of-window tokens to get in. Compressing key states before applying RoPE, FreqKV extends the context window without requiring position extrapolation or interpolation.
(a) GemFilter.
(b) FastKV.
(c) SnapKV.
(d) ThinK.
(e) PyramidKV.
(f) FreqKV.
@@ 4.3 Long Context Understanding
Figure 4: The Needle-in-a-Haystack results of different KV compression methods on LLaMA-3-instruct-8B, with the x-axis representing the document length ranging from 1K to 16K tokens, and the y-axis showing the position of the “needle” within the document.
Needle-in-a-Haystack.
Figure 4 illustrates evaluation results on Needle-in-a-Haystack. FreqKV performs well in extending the context window of LLaMA-3-instruct. In contrast, other KV compression methods fail beyond the original window boundary. This breakdown is consistent with our observations in RULER.
4.4 Long Context Generation
We use LongGenBench to evaluate the long-context generation performance of FreqKV. This benchmark constructs extended generation tasks by concatenating multiple chain-of-thought questions and answers, followed by several real questions in GSM8K and MMLU for long-context responses. The model is then required to generate answers with a maximum output length of 4096 tokens.
Table 5: Performance on LongGenBench.
Method GSM8K MMLU
LLaMA-3 42.8 23.8
+D2O 8.0 24.8
+GemFilter 48.5 25.4
+FastKV 48.5 25.4
+SnapKV 49.6 25.8
+ThinK 45.5 25.7
+PyramidKV 49.6 26.2
  +FreqKV 54.2 31.2
The accuracy scores (%) are provided in Table 5. FreqKV demonstrates substantial improvements in long-generation capability and consistently outperforms existing KV compression methods across both GSM8K and MMLU. FreqKV preserves essential contextual information through iterative frequency-domain compression, enabling robust long-context decoding.
@@ 5 Analysis
5.1 Efficient Extrapolation
Table 6: Extrapolation performance of FreqKV on PG-19 up to 256K context length. We measure latencies (seconds) and PPL scores at different lengths. “Compression Count” denotes the number of compressions during decoding. “Compression Latency” stands for the latency introduced by all the frequency-domain compressions during decoding.
Evaluation Length 8K 16K 32K 64K 128K 256K
Compression Count 3 7 15 31 63 127
Decoding Latency 214.96 445.69 907.49 1832.77 3692.03 7422.64
Compression Latency 0.62 (0.29%) 1.09 (0.24%) 2.18 (0.24%) 4.00 (0.22%) 8.31 (0.23%) 16.39 (0.22%)
PPL 7.76 7.74 7.73 7.73 7.73 7.73
Furthermore, we evaluate the performance of FreqKV at longer lengths on the validation set of PG-19 as shown in Table 6. With the training length limited to 8K tokens, it effectively extrapolates the context window of LLaMA-2-7B to 256K, which is 64 times larger than the original size. This is enabled by its iterative compression strategy, which applies fewer compressions to recent tokens while progressively compressing earlier ones. Consequently, FreqKV maintains low perplexity as the evaluation context length scales from 8K to 256K, far exceeding the training length.
In addition, latencies (seconds) of decoding and frequency-domain compression at different lengths are also reported in Table 6. It shows that the decoding time of FreqKV increases approximately linearly, with a negligible time spent on compression. The overhead of the compression accounts for less than 0.3% and does not grow with the context length since the number of compressions increases linearly with the context length. Further analysis for the compression overhead is provided in Appendix G. In summary, FreqKV demonstrates strong extrapolation ability to ultra-long contexts while preserving both efficiency and stability.
5.2 Training Cost
The training time and memory requirements of FreqKV and Full FT on LLaMA-2-7B are reported in Table 7. 8 NVIDIA RTX6000 Ada-48GB GPUs are used. For comparison, we set the batch size per device to 1 and gradient accumulation steps to 8. Statistics show that FreqKV significantly reduces both training time (by 40.89%) and memory requirements (by 44.19%), demonstrating its computational efficiency.
Method Training Length Time (hours) Memory (GB)
Full FT 8K 28.88 44.04
FreqKV 8K 17.07 24.58
16K 38.57 34.00
32K 89.35 45.37